### v1

# 360VOTS: Visual Object Tracking and Segmentation in Omnidirectional Videos

Abstract
Visual object tracking and segmentation in omnidirectional videos are challenging due to the wide field-of-view and large spherical distortion brought by 360°$\degree$ images. To alleviate these problems, we introduce a novel representation, extended bounding field-of-view (eBFoV), for target localization and use it as the foundation of a general 360 tracking framework which is applicable for both omnidirectional visual object tracking and segmentation tasks. Building upon our previous work on omnidirectional visual object tracking (360VOT), we propose a comprehensive dataset and benchmark that incorporates a new component called omnidirectional video object segmentation (360VOS). The 360VOS dataset includes 290 sequences accompanied by dense pixel-wise masks and covers a broader range of target categories. To support both the development and evaluation of algorithms in this domain, we divide the dataset into a training subset with 170 sequences and a testing subset with 120 sequences. Furthermore, we tailor evaluation metrics for both omnidirectional tracking and segmentation to ensure rigorous assessment. Through extensive experiments, we benchmark state-of-the-art approaches and demonstrate the effectiveness of our proposed 360 tracking framework and training dataset.
Index Terms:
Dataset, omnidirectional vision, visual object tracking, video object segmentation
I Introduction
Visual object tracking (VOT) and video object segmentation (VOS) are two essential tasks in computer vision for spatially and temporally localizing target objects in video sequences. A robust and precise approach is demanded in various applications such as video analysis, human-machine interaction, and intelligent robots. While both VOT and VOS are integral to understanding object dynamics in videos, their fundamental differences are in the level of localization granularity. VOT approaches typically utilize the bounding box to represent the target position and rough size. In contrast, VOS approaches seek to provide pixel-wise localization of the object’s silhouette frame-by-frame. In the last decade, a remarkable development can be seen in these tasks with the flourish of excellent trackers like [1, 2, 3, 4, 5] in VOT and [6, 7, 8, 9, 10] in VOS, as well as various benchmarks including [11, 12, 13, 14, 15] for VOT and [16, 17, 18, 19] for VOS. Whereas most existing research focuses on perspective scenarios, there is little attention paid to object tracking and segmentation in the omnidirectional view.
Omnidirectional object tracking and segmentation employs a 360°$\degree$ camera to track the target object. With its omnidirectional field-of-view (FoV), a 360°$\degree$ camera offers continuous observation of the target over a longer period, minimizing the out-of-view issue. This advantage is crucial for intelligent agents to achieve stable, long-term tracking and perception. In general, an ideal spherical camera model is used to describe the projection relationship of a 360°$\degree$ camera. The 360°$\degree$ image is widely represented by equirectangular projection (ERP) [20], which has two main features: 1) crossing the image border and 2) extreme distortion as the latitude increases. Moreover, due to inherent limitations or manufacturing defects of the camera, the 360°$\degree$ image may suffer from stitching artifacts that would blur, break, or duplicate the shape of objects. Meanwhile, omnidirectional FoV means it is inevitable to capture the photographers. They would distract and occlude the targets. These phenomena are illustrated in Figure 1. Eventually, they bring new challenges to perform object tracking and segmentation in 360°$\degree$ videos.
To tackle these challenges, we first explore a new representation in omnidirectional tracking that utilizes extended bounding field-of-view (eBFoV) for precise target localization. Compared to the commonly used bounding box (BBox), BFoV  [21, 22] represents object localization on the unit sphere in an angular fashion. This representation takes the spherical nature of 360°$\degree$ images into account, providing improved constraints for accurately localizing objects and eliminating dependence on image resolution. We therefore can utilize BFoV to extract less-distorted images for target searching. To enable proper handling of objects even when they occupy significant portions of the omnidirectional image, we further extend the BFoV definition and propose eBFoV to accommodate the extraction of search regions. By leveraging the benefits of eBFoV, we propose a general 360 tracking framework that effectively enhances the performance of the conventional local trackers devised for perspective scenes in omnidirectional scenes.
Furthermore, we present 360VOTS, a challenging benchmark dataset to facilitate research and evaluation on omnidirectional tracking and segmentation, including 360VOT [23] and 360VOS datasets. 360VOT provides dense (rotated) BBox and BFoV annotations of 120 test sequences. Inheriting parts of sequences from the 360VOT dataset, 360VOS incorporates additional 191 sequences and 30 tracking object categories. This results in a large-scale dataset consisting of a total of 290 sequences in 62 categories. Importantly, 360VOS provides dense pixel-wise annotations as ground truth and assigns 170 sequences as the training set. These pixel-wise annotations are utilized to obtain unbiased (rotated) BBoxes and BFoVs ground truth, ensuring comprehensive evaluation and comparison of different approaches. To accurately evaluate the performance of state-of-the-art (SOTA) approaches on omnidirectional tracking and segmentation, we develop new metrics tailored for 360°$\degree$ images. These metrics encompass dual success, dual (normalized) precision, and angle precision to assess tracking performance, as well as spherical region similarity and spherical contour accuracy to evaluate segmentation performance. Moreover, we exploit the 360VOS training set to retrain the tracker and apply our 360 tracking framework to achieve new SOTA results. With the introduction of new representations, datasets, metrics, and baselines, we believe 360VOTS serves as a vital resource to foster development in omnidirectional visual object tracking and segmentation.
In summary, the contribution of this work includes:
  • We explore the new representations for object location in 360°$\degree$ images, as well as a general 360 tracking framework.
  • Our proposed 360VOTS is the first benchmark dataset for both omnidirectional visual object tracking and segmentation. We provide densely annotated masks for segmentation and subsequently convert masks into four types of unbiased ground truth for tracking.
  • We propose a set of new metrics specifically designed to enable rigorous performance evaluation of omnidirectional tracking and segmentation algorithms.
  • We benchmark 20 tracking algorithms and 16 segmentation algorithms on 360VOT and 360VOS respectively, and develop new baselines for future comparisons.
Figure 1: Example sequences and annotations of 360VOTS benchmark dataset. The target objects in each 360°$\degree$ frame are annotated with masks which can be converted into four different representations as ground truth, including bounding box (BBox), rotated bounding box (rBBox), bounding field-of-view (BFoV), and rotated bounding field-of-view (rBFoV). 360VOTS brings distinct challenges for tracking and segmentation, i.e., crossing border (CB), large distortion (LD), and stitching artifact (SA).
TABLE I: Comparison of popular visual object tracking (VOT) and video object segmentation (VOS) benchmarks.
Benchmark Videos test / train Total frames Min frames Mean frames Median frames Max frames Object classes Attr. Annotation Feature Year
ALOV300[24] 314 152K 19 483 276 5,975 64 14 sparse BBox diverse scenes 2013
OTB100[11] 100 81K 71 590 393 3,872 16 11 dense BBox short-term 2015
NUS-PRO[25] 365 135K 146 371 300 5,040 8 12 dense BBox occlusion-level 2015
TC128[26] 129 55K 71 429 365 3,872 27 11 dense BBox color enhanced 2015
UAV123[12] 123 113K 109 915 882 3,085 9 12 dense BBox UAV 2016
DTB70[27] 70 16K 68 225 202 699 29 11 dense BBox UAV 2016
NfS[28] 100 383K 169 3,830 2,448 20,665 17 9 dense BBox high FPS 2017
UAVDT[29] 100 78K 82 778 602 2,969 27 14 sparse BBox UAV 2017
OxUvA[30] 337 1.55M 900 4,260 2,628 37,440 22 6 sparse BBox long-term 2018
TrackingNet[31] 511/30K 226K/14M 96/87 441/471 390/480 2,368/690 27 15 sparse BBox large scale 2018
LaSOT[14] 280/1,120 685K/2,816K 1,000/898 2,448/2,514 2,102/2,023 9,999/11K 85 14 dense BBox category balance 2018
GOT-10k[15] 420/9,335 56K/1,403K 29/29 127/150 100/101 920/1,481 84 6 dense BBox generic 2019
TOTB[32] 225 86K 126 381 389 500 15 12 dense BBox transparent 2021
TREK-150[33] 150 97K 161 649 484 4,640 34 17 dense BBox FPV 2021
VOT[13] 62 20K 41 321 242 1,500 37 9 dense BBox annual 2022
360VOT[23] 120 113K 251 940 775 2,400 32 20 dense (r)BBox & (r)BFoV 360°$\degree$ images 2023
FBMS-59[16] 30/29 7,306/6,554 19/19 244/226 210/193 800/720 - - sparse Mask moving objects 2014
DAVIS[18] 30/90 2,860/6,208 31/25 70/69 70/70 127/104 376 15 dense Mask dense annotated 2017
YouTube-VOS[19] 508/3471 70K/470K 36/20 139/135 150/150 180/180 94 - sparse Mask comprehensive 2018
LVOS[17] 50/120 30K/65K 257/200 598/544 489/414 2,000/2,280 27 13 dense Mask long-term 2023
360VOS 120/170 98K/144K 111/62 821/845 684/714 2,400/2,402 62 20 dense Mask 360°$\degree$ images 2024
II Related Work
II-A Benchmarks for Visual Object Tracking
With the remarkable development of the visual object tracking community, previous works have proposed numerous benchmarks in various scenarios.
ALOV300 [24] is a sparse benchmark introducing 152K frames and 16K annotations, while UAVDT [29] focuses on UAV scenarios and has 100 videos. TrackingNet [31] is a large-scale dataset collecting more than 14M frames based on the YT-BB dataset [34]. As YT-BB only provides fine-grained annotations at 1 FPS, they explored a tracker to densify the annotations without further manual refinement. OxUvA [30] targets long-term tracking by constructing 337 video sequences, but each video only has 30 frames annotated.
One of the first dense BBox benchmarks is OTB100 [11] which is extended from OTB50 [35] and has 100 sequences. NUS-PRO [25] takes the feature of occlusion-level annotation and provides 365 sequences, while TC128 [26] researches the chromatic information in visual tracking. UAV123 [12] and DTB70 [27] offer 123 and 70 aerial videos of rigid objects and humans in various scenes. NfS [28] consists of more than 380K frames captured at 240 FPS studying higher frame rate tracking, while LaSOT [14] is a large-scale and category balanced benchmark of premium quality. GOT-10k [15] provides about 1.5M annotations and 84 classes of objects, aiming at generic object tracking. The annual tracking challenge VOT [13] offered 62 sequences and 20K frames in 2022. A more recent benchmark TOTB [32] mainly focuses on transparent object tracking. TREK-150 [33] introduces 150 sequences of tracking in First Person Vision (FPV) with the interaction between the person and the target object.
Differently, our 360VOT focuses on object tracking and explores new representations on omnidirectional videos. A summarized comparison with existing VOT benchmarks is reported in the upper part of Table I.
II-B Benchmarks for Video Object Segmentation
There are several benchmarks for video object segmentation proposed in the last decade as described in Table I. FBMS-59[16] acknowledges the importance of motion information and concentrates on motion segmentation, providing a total of 13,860 frames across 59 sequences. DAVIS[18] is the first dataset designed for video object segmentation (VOS), annotating 376 classes of objects on each frame of 120 short videos. Leveraging the diverse scenarios and photography conditions on YouTube videos, YouTube-VOS[19] introduces a large-scale VOS dataset with more than 540K frames and contains 3471 and 508 sequences as the training set and the test set separately. Each video clip has a duration of 3 to 6 seconds and is annotated every 5 frames. One of the most recent datasets is LVOS[17] which is the first long-term VOS benchmark dataset with an average length of 560 densely labeled frames. In contrast, our 360VOS provides pixel-wise masks of omnidirectional images for both training and testing.
II-C Adapting Object Tracking Schemes in Segmentation
To guarantee high tracking speed, the trackers for single object tracking generally crop the image and search for the target in small local regions. The tracking scheme is vital in selecting searching regions and interpreting network predictions over sequences in the inference phase. A compatible tracking inference scheme can enhance tracking performance. For example, DaSiamRPN [36] explored a local-to-global searching strategy for long-term tracking. SiamX [37] proposed an adaptive inference scheme to prevent tracking loss and realized fast target re-localization. Similarly, localizing the target objects is also vital for video object segmentation tasks. Both RTS[38] and SiamMask[39] demonstrated that pixel-wise masks can provide a more accurate estimation of bounding boxes for VOT tasks. Additionally, their tracking schemes improve the robustness of VOS methods and contribute to target localization in long-term videos.
In 360VOTS, we take advantage of eBFoV to localize the targets on 360°$\degree$ videos and obtain less-distorted search regions. It allows us to exploit arbitrary local visual trackers, which are trained on normal perspective images to achieve enhanced performance on 360°$\degree$ object tracking and segmentation.
III Tracking on 360°$\degree$ Video
The 360°$\degree$ video is composed of frames using the most common ERP. Each frame can capture 360°$\degree$ horizontal and 180°$\degree$ vertical field of view. Although omnidirectional FoV avoids out-of-view issues, the target may cross the left and right borders of a 2D image. Additionally, nonlinear projection distortion makes the target largely distorted when they are near the top or bottom of the image, as illustrated in Figure 1. Therefore, a new representation and framework that fits the 360°$\degree$ video for omnidirectional visual tracking is necessary.
Figure 2: Pickup Truck. The comparison of the bounding regions of different representations on a 360°$\degree$ image. The unwarped images based on BFoV and rBFoV are less distorted.
Figure 3: The boundaries on the 360°$\degree$ images and the corresponding unwarped images of different BFoV definitions. The tangent BFoV is displayed in blue and the extended BFoV is in orange. $M$ on the sphere surface denotes the object center and tangent point. Blue plane with dotted borders represents a larger plane out of space. Best viewed in color.
III-A Representation for the Target Location
The (r)BBox is the most common and simple way to represent the target object’s position in perspective images. It is a rectangular area defined by the rectangle around the target object on the image and denoted as $[cx,cy,w,h,\gamma]$, where $cx,cy$ are object center, $w,h$ are width and height. The rotation angle $\gamma$ of BBox is always zero. However, these representations would become less accurate to properly constrain the target on the 360°$\degree$ image, as depicted in Figure 2. The works [40, 22] for 360°$\degree$ object detection show that the BFoV and rBFoV are more appropriate representations on 360°$\degree$ images. Basically, we can use the spherical camera model $\mathcal{P}$ to formulate the mathematical relationship between the 2D image in ERP and a continuous 3D unit sphere [41]. (r)BFoV is then defined as $[clon,clat,\theta,\phi,\gamma]$ where $[clon,clat]$ are the longitude and latitude coordinates of the object center at the spherical coordinate system, $\theta$ and $\phi$ denote the maximum horizontal and vertical field-of-view angles of the object’s occupation. The represented region of (r)BFoV on the 360°$\degree$ image is commonly calculated via a tangent plane [40, 22], $T(\theta,\phi)\in\mathbb{R}^{3}$, and formulated as:
$I(\mbox{\footnotesize(r)BFoV}\,|\,\Omega)=\mathcal{P}(\mathcal{R}_{y}(clon)\cdot\mathcal{R}_{x}(clat)\cdot\mathcal{R}_{z}(\gamma)\cdot\Omega),$ (1)
where,
$\small\mathcal{R}_{y}(clon)$ $=\begin{bmatrix}cos(clon)&0&sin(clon)\\ 0&1&0\\ -sin(clon)&0&cos(clon)\end{bmatrix},$ (2)
$\mathcal{R}_{x}(clat)$ $=\begin{bmatrix}1&0&0\\ 0&cos(clat)&-sin(clat)\\ 0&sin(clat)&cos(clat)\end{bmatrix},$ (3)
$\mathcal{R}_{z}(\gamma)$ $=\begin{bmatrix}cos\gamma&-sin\gamma&0\\ sin\gamma&cos\gamma&0\\ 0&0&1\end{bmatrix},$ (4)
$\Omega=T$ $=\begin{bmatrix}\mathbf{X}\\ \mathbf{Y}\\ \mathbf{Z}\end{bmatrix}=\begin{bmatrix}-tan(\theta/2):tan(\theta/2)\\ -tan(\phi/2):tan(\phi/2)\\ 1\end{bmatrix}.$ (5)
The projection function $\mathcal{P}$ to depict the relationship between the 3D camera space $[X,Y,Z]$ and the 2D image space $[u,v]$ is formulated as:
$u$ $=(\frac{lon}{2\pi}+0.5)*W=arctan(X/Z),$ (6)
$v$ $=(-\frac{lat}{\pi}+0.5)*H=arctan(\frac{-Y}{\sqrt{X^{2}+Z^{2}}}),$
where $-\pi<lon<\pi$ and $-\pi/2<lat<\pi/2$ denote the longitude and latitude in the spherical coordinate system. W and H are the width and height of the 360°$\degree$ image. The unwarped images based on tangent BFoV are distortion-free under the small FoV, as shown in Figure 2.
However, this definition has a disadvantage on large FoV and cannot represent the region exceeding 180°$\degree$ FoV essentially. With the increasing FoV, the unwarped images from the tangent planes have drastic distortions, shown in the upper row in Figure 3. This defect limits the application of BFoV on visual object tracking since trackers rely on unwarped images for target searching. To address this problem, we extended the definition of BFoV. When the bounding region involves a large FoV, i.e., larger than 90°$\degree$, the extended BFoV leverages a spherical surface $S(\theta,\phi)\in\mathbb{R}^{3}$ instead of a tangent plane to represent the bounding region on the 360°$\degree$ image:
$S=\begin{bmatrix}cos(\Phi)sin(\Theta)&-sin(\Phi)&cos(\Phi)cos(\Theta)\end{bmatrix}^{T},\$ (7)
where $\Phi\in[-\phi/2,\phi/2],\Theta\in[-\theta/2,\theta/2]$. Therefore, the corresponding region of extended (r)BFoV on 360°$\degree$ image is
$I(\mbox{\footnotesize(r)BFoV}\,|\,\Omega),\quad\Omega=\begin{cases}T(\theta,\phi),&\theta<90\degree,\phi<90\degree\\ S(\theta,\phi),&otherwise\end{cases}.$ (8)
The comparisons of the boundary on 360°$\degree$ images and corresponding unwarped images based on tangent BFoV and the extended BFoV are shown in Figure 3. For simplicity, BFoV in the following content denotes extended BFoV unless explicitly stated otherwise.
III-B 360 Tracking Framework
To conduct omnidirectional tracking and segmentation using an existing local VOT or VOS tracker, we propose a 360 tracking framework, as depicted in Figure 4. The framework leverages extended BFoV to address challenges caused by object crossing-border and large distortion on the 360°$\degree$ image.
Given an initial BFoV or mask, the framework first calculates the corresponding search region $I$ on the 360°$\degree$ image via Eq. 8. Subsequently, by remapping ($\tau$) the 360°$\degree$ image using corresponding pixel coordinates recorded in $I$, it extracts a less distorted local image for target identification. Within the extracted image, a local visual tracker then infers a (r)BBox or a segmentation mask. Finally, we utilize $I$ to convert the local prediction back to the global bounding region on the 360°$\degree$ image via inverse projection $\tau^{-1}$. For (r)BBox prediction, we calculate the minimum area (rotated) rectangle on the 360°$\degree$ image. For (r)BFoV, we re-project the bounding region’s coordinates onto the spherical coordinate system and calculate the maximum bounding FoV. As for the final mask in 360°$\degree$ images, we dilate the reprojected mask to avoid erosion artifacts resulting from inverse projection. To maintain continuous tracking, the BFoV search region for the subsequent frame is updated by the current predicted BFoV or mask. In this manner, the search region follows the motions of the target throughout the entire sequence to achieve object tracking and segmentation. It is noted that the framework does not rely on nor affect the network architecture of the tracker, allowing for the adoption of an arbitrary local visual tracker for omnidirectional tracking and segmentation.
Figure 4: The 360 tracking framework. The framework focuses on extracting a less-distorted search region to enable precise local visual tracking. The general framework accommodates arbitrary local visual trackers to generate masks or 4 types of representation. $\tau$ is the remapping operation to extract the less-distorted search region from the input 360°$\degree$ image, while $\tau^{-1}$ is its inverse operation.
IV Benchmark Dataset: 360VOTS
IV-A Collection and Annotation
The resources of 360°$\degree$ videos in 360VOTS were collected from YouTube or captured using a 360°$\degree$ camera. We ranked and filtered hundreds of candidate videos considering tracking difficulty scale and some additional challenging cases. Videos with additional challenges were assigned a higher priority (i.e. distinguishing targets from other highly comparable objects). Then, videos were further selected and sampled into sequences with a frame number threshold ($\leq$2400). The relatively stationary frames were further discarded manually. All selected frames were annotated by an interactive segmentation tool [23]. Considering distribution balance, 120 sequences are finally selected as the 360VOT benchmark and 290 sequences as the 360VOS dataset, sharing 99 sequences for tracking and segmentation benchmarks.
TABLE II: Attribute description. The 360VOTS contains 13 basic attributes widely used in the existing benchmarks, and 7 distinct challenging attributes as described in the last block of the row.
Attr. Meaning
IV The target is subject to light variation.
BC The background has a similar appearance as the target.
DEF The target deforms during tracking.
MB The target is blurred due to motion.
CM The camera has abrupt motion.
ROT The target rotates related to the frames.
POC The target is partially occluded.
FOC The target is fully occluded.
ARC The ratio of the annotation aspect ratio of the first and the current frame is outside the range [0.5, 2].
SV The ratio of the annotation area of the first and the current frame is outside the range [0.5, 2].
FM The motion of the target center between contiguous frames exceeds its own size.
LR The area of the target annotation is less than $1000$ pixels.
HR The area of the target annotation is larger than $500^{2}$ pixels.
SA The 360°$\degree$ images have stitching artifacts and they affect the target object.
CB The target is crossing the border of the frame and partially appears on the other side.
FMS The motion angle on the spherical surface of the target center is larger than the last BFoV.
LFoV The vertical or horizontal FoV of the BFoV is larger than $90\degree$.
LV The range of the latitude of the target center across the video is larger than $50\degree$.
HL The latitude of the target center is outside the range $[-60\degree,60\degree]$, lying in the “frigid zone”.
LD The target suffers large distortion due to the equirectangular projection.
IV-B Attribute
Each sequence is annotated with a total of 20 different attributes: illumination variation (IV), background clutter (BC), deformable target (DEF), motion blur (MB), camera motion (CM), rotation (ROT), partial occlusion (POC), full occlusion (FOC), aspect ratio change (ARC), scale variation (SV), fast motion (FM), low resolution (LR), high resolution (HR), stitching artifact (SA), crossing border (CB), fast motion on the sphere (FMS), large FoV (LFoV), latitude variation (LV), high latitude (HL) and large distortion (LD).
The detailed meaning of each attribute is described in Table II. Among them, IV, BC, DEF, MB, CM, ROT, POC, and LD attributes are manually labeled, while the others are computed from the annotation results of targets. The distinct features of the 360°$\degree$ image are well represented in 360VOTS: location variations (FMS, LFoV, and LV), external disturbances (SA and LD) and special imaging (CB and HL).
Figure 5 illustrates the comparison between 360VOT and 360VOS in terms of sequence counts and dependencies for each attribute. As shown in Figure 5a, most attributes have similar proportions in 360VOT and 360VOS. In 360VOT, the scale changes (ARC and SV) and motion (MB and FM) are common challenges illustrated in Figure 5b. By contrast, in 360VOS, the scale changes (ARC and SV) remain the predominant features, whereas camera motion (CM) and target rotation (ROT) gain an obvious increase in the number of sequences illustrated in Figure 5c.
(a) Attribute histograms.
(b) 360VOT dependency wheel.
(c) 360VOS dependency wheel.
Figure 5: Attribute distributions and dependencies in 360VOTS. The additional introduced attributes are highlighted in red palettes, while the commonly used attributes in existing works are highlighted in blue palettes. The arc lengths on the dependency wheel measure the dominance of the attributes in the dataset. Thicker connection lines indicate a higher frequency of co-presence between attribute pairs, and vice versa.
(a) The category tree of 360VOT.
(b) The category tree of 360VOS.
Figure 6: The tree maps of the category structures for 360VOTS. Newly introduced classes in 360VOS are highlighted in red, while the classes that have been re-categorized from the original 360VOT are highlighted in blue.
IV-C Dataset Statistics of 360VOTS
360VOTS cover four primary categories of humans, animals, rigid objects, and human & carrier cases. As shown in Figure 6, the primary classes are depicted on the branches, with the leaves denoting the respective sub-classes within each primary class. In the majority of sequences, a single target is annotated across all frames, while up to two components of a single target (e.g. bicyclist and bicycle) are annotated separately in several sequences under the human & carrier cases class. Overall, the 120 sequences in the 360VOT dataset are grouped into 32 sub-classes and used for evaluation only. The 360VOS dataset has 62 sub-classes and covers more expansive sequences. In addition, we split the 360VOS dataset into training and test sets with balanced classes of tracking targets, avoiding class bias between the two sets. Eventually, the test set of 360VOS comprises 98,487 annotated frames across 120 sequences and the training set includes 143,634 annotated frames within 170 sequences. As a benchmark suitable for long-term tracking and segmentation evaluation, 360VOS is characterized by an average of 821 frames per sequence in the test set and 845 frames per sequence in the training set, respectively, as reported in Table I.
V Experiments
V-A Evaluation Metrics
Evaluation on bounding regions. To conduct the experiments on the tracking results, we use the standard one-pass evaluation (OPE) protocol [11] and measure the success S, precision P, and normalized precision $\overline{P}$ [31] of the trackers over the test sequences. Success S is computed as the intersection over union (IoU) between the tracking results $B^{tr}$ and the ground truth annotations $B^{gt}$, where the $B^{tr/gt}$ indicates a (r)BBox. The trackers are ranked by the area under curve (AUC), which is the average of the success rates corresponding to the sampled thresholds $[0,1]$. The precision P is computed as the distance between the results $\mathbf{C}^{tr}$ and the ground truth centers $\mathbf{C}^{gt}$. The trackers are ranked by the precision rate on the specific threshold (i.e., 20 pixels). The normalized precision $\overline{\mbox{P}}$ is scale-invariant, which normalizes the precision P over the size of the ground truth and then ranks the trackers using the AUC for the $\overline{\mbox{P}}$ between 0 and 0.5. For the perspective image using (r)BBox, these metrics can be formulated as:
$\begin{split}\mbox{S}=IoU(B^{gt},B^{tr}),\,\mbox{P}=||\mathbf{C}^{gt}_{xy}-\mathbf{C}^{tr}_{xy}||_{2}\\ \overline{\mbox{P}}=||diag(B^{gt},B^{tr})(\mathbf{C}^{gt}_{xy}-\mathbf{C}^{tr}_{xy})||_{2}.\end{split}$ (9)
However, for 360°$\degree$ images, the target predictions may cross the image. To handle this situation and increase the accuracy of BBox evaluation, we introduce dual success Sdual and dual precision Pdual. Specifically, we shift the $B^{gt}$ to the left and right by $W$, the width of 360°$\degree$ images, to obtain two temporary ground truth $B^{gt}_{l}$ and $B^{gt}_{r}$. Based on the new ground truth, we then calculate extra success Sl and Sr and precision Pl and Pr using Eq. 9. Finally, Sdual and Pdual are measured by:
$\mbox{S}_{dual}$ $=max\{\mbox{S}_{l},\mbox{S},\mbox{S}_{r}\},$ $\mbox{P}_{dual}$ $=min\{\mbox{P}_{l},\mbox{P},\mbox{P}_{r}\}.$ (10)
Sdual and S, as Pdual and P, are the same when the annotation does not cross the image border. Similarly, we can compute the dual normalized precision $\overline{\mbox{P}}_{dual}$.
Since objects suffer significant non-linear distortion in the polar regions due to the equirectangular projection, the distance between the predicted and ground truth centers may be large on the 2D image but they are adjacent on the spherical surface. It means that dual precision Pdual is sensitive for 360°$\degree$ images. Therefore, we propose a new metric Pangle, which is measured as the angle precision between the vectors of the ground truth and the tracker results in the spherical coordinate system. The angle precision Pangle is formulated as:
$\mbox{P}_{angle}=||\mathbf{C}^{gt}_{lonlat}-\mathbf{C}^{tr}_{lonlat}||_{2}\,.$ (11)
The different trackers are ranked with angle precision rate on a threshold, i.e., 3°$\degree$. Moreover, when target positions are represented by BFoV or rBFoV, we utilize spherical IoU [21] to compute the success metric, denoted as spherical success Ssphere, while only Ssphere and Pangle are measured.
Evaluation on segmentation. Region similarity $\mathcal{J}$ and contour accuracy $\mathcal{F}$ [42] are two common metrics in evaluation of segmentation results. Region similarity $\mathcal{J}$ measures the pixel-based IoU between an output segmentation $M^{tr}$ and its ground truth mask $M^{gt}$. $\mathcal{F}$ computes the F1-score by precision and recall on contour $\mbox{P}^{c}$ and $\mbox{R}^{c}$ between the contours of $M^{tr}$ and $M^{gt}$. The two standard metrics are defined as:
$\small\mathcal{J}$ $=IoU(M^{gt},M^{tr}),\!\!\!\!\!\!\!\!\!\!\!\!$$\mathcal{F}$ $=\frac{2\mbox{P}^{c}\mbox{R}^{c}}{\mbox{P}^{c}+\mbox{R}^{c}}.$ (12)
When a target is situated in high-latitude areas (i.e., close to the top or bottom) on 360°$\degree$ images, its segmentation is usually stretched due to the spherical distortion, leading to issues of overestimation in the metrics of $\mathcal{J}$ and $\mathcal{F}$. Figure 7a shows a sample target of the mentioned case, as well as its mask and pixel grids in the perspective view (i.e., parallel with the 360°$\degree$ image boundaries) are marked separately. To handle this issue, we re-project the 360°$\degree$ image on a 3D sphere as shown in Figure 7b. The spherical pixel grids near the polar regions contribute less weight to the segmentation than other regions. Considering the weights varying along the latitude, we apply pre-computed spherical weights $\mathfrak{W}$ on the standard metrics $\mathcal{J}$ and $\mathcal{F}$ obtaining two spherical metrics, i.e., spherical region similarity $\mathcal{J}_{sphere}$ and spherical contour accuracy $\mathcal{F}_{sphere}$.
Specifically, in a 360°$\degree$ image with width $W$ and height $H$, the spherical weight measuring the importance of a specific pixel $[u,v]$ can be defined by the corresponding surface area on a unit sphere (radius $r=1$), formulated as:
$\mathfrak{W}_{u,v}=\|\int_{\varphi-\Delta\varphi/2}^{\varphi+\Delta\varphi/2}\int_{\vartheta-\Delta\vartheta/2}^{\vartheta+\Delta\vartheta/2}\sin\vartheta\;d\vartheta d\varphi\|,$ (13)
where $[\vartheta,\varphi]$ is the longitude and latitude of the spherical coordinate which can be computed from $[u,v]$ via Eq. 6, $\Delta\vartheta=\pi/H$ and $\Delta\varphi=2\pi/W$.
With the pre-computed spherical weights, spherical region similarity $\mathcal{J}_{sphere}$ is defined by:
$\mathcal{J}_{sphere}=IoU(M^{gt}\odot\mathfrak{W},M^{tr}\odot\mathfrak{W}).$ (14)
For evaluating contour accuracy on a sphere $\mathcal{F}_{sphere}$, we adopt the bipartite graph matching [43] as in [42]. The spherical contour accuracy $\mathcal{F}_{sphere}$ is measured by:
$\begin{split}\mbox{P}_{sphere}^{c}&=\frac{\sum\Psi(c(M^{tr})|c(M^{gt}))\odot\mathfrak{W}}{\sum c(M^{tr})\odot\mathfrak{W}},\\ \mbox{R}_{sphere}^{c}&=\frac{\sum\Psi(c(M^{gt})|c(M^{tr}))\odot\mathfrak{W}}{\sum c(M^{gt})\odot\mathfrak{W}},\\ \mathcal{F}_{sphere}&=\frac{2\mbox{P}_{sphere}^{c}\mbox{R}_{sphere}^{c}}{\mbox{P}_{sphere}^{c}+\mbox{R}_{sphere}^{c}},\end{split}$ (15)
where the matching operation $\Psi(\cdot|\cdot)$, the contour of a mask $c(\cdot)$, and the element-wise multiplication $\odot$.
(a) Sample 360°$\degree$ image with a target elephant and its corresponding mask with perspective pixel-grids.
(b) Projection of the mask and grids on a spherical surface, where the size of the grids varies with latitude.
Figure 7: Difference between perspective pixel grids on a 360°$\degree$ image and spherical pixel grids on a 3D sphere.
V-B VOT Evaluation
VOT baseline trackers. We evaluated 20 state-of-the-art visual object trackers on 360VOT, as shown in Table III. According to the latest development of visual tracking, the compared methods can be roughly classified into three groups: transformer trackers, Siamese trackers, and other deep learning-based trackers. Specifically, the transformer trackers contain Stark [44], ToMP [45], MixFormer [46], SimTrack [47] and AiATrack [48]. The Siamese trackers include SiamDW [49], SiamMask [39], SiamRPNpp [4], SiamBAN [50], AutoMatch [51], Ocean [52] and SiamX [37]. For other deep trackers, UDT [53], Meta-SDNet [54], MDNet [2], ECO [3], ATOM [55], KYS [56], DiMP [5], PrDiMP [57] were evaluated. We used the official implementation, trained models, and default configurations to ensure a fair comparison among trackers. In addition, we developed a new baseline AiATrack-360 that combines the transformer tracker AiATrack [48] with our 360 tracking framework. We also adapted a different kind of tracker SiamX [37] with our framework, named SiamX-360, to verify the generality of the proposed framework.
TABLE III: Overall performance on 360VOT BBox in terms of dual success, dual precision, dual normalized precision, and angle precision.
Bold blue indicates the best results in the tracker group. Bold red indicates the best results overall.
VOT Tracker 360VOT BBox
$\mbox{S}_{dual}$ (AUC) $\mbox{P}_{dual}$ $\overline{\mbox{P}}_{dual}$ (AUC) $\mbox{P}_{angle}$
UDT[53] 0.104 0.075 0.117 0.098
Meta-SDNet[54] 0.131 0.097 0.164 0.136
MDNet[2] 0.150 0.106 0.188 0.143
ECO[3] 0.175 0.130 0.212 0.179
ATOM[55] 0.252 0.216 0.286 0.266
KYS[56] 0.286 0.245 0.312 0.296
DiMP[5] 0.290 0.247 0.315 0.299
PrDiMP[57] 0.341 0.292 0.371 0.347
SiamDW[49] 0.156 0.116 0.190 0.156
SiamMask[39] 0.189 0.161 0.220 0.203
SiamRPNpp[4] 0.201 0.175 0.233 0.213
SiamBAN[50] 0.205 0.187 0.242 0.227
AutoMatch[51] 0.208 0.202 0.261 0.248
Ocean[52] 0.240 0.223 0.287 0.264
SiamX[37] 0.302 0.265 0.331 0.315
Stark[44] 0.381 0.356 0.403 0.408
ToMP[45] 0.393 0.352 0.421 0.413
MixFormer[46] 0.395 0.378 0.417 0.424
SimTrack[47] 0.400 0.373 0.421 0.424
AiATrack[48] 0.405 0.369 0.427 0.423
SiamX-360 0.391 0.365 0.430 0.425
AiATrack-360 0.534 0.506 0.563 0.574
(d) (a) Crossing Border
(e) (b) Fast Motion on the Sphere
(f) (c) Latitude Variation
Figure 8: Comparing BBox tracking performances in 360VOT of different trackers, in terms of dual success rate and angle precision rate under the three distinct attributes of 360VOT.
TABLE IV: Tracking performance based on other annotations of 360VOT using 360 tracking framework.
VOT Tracker 360VOT rBBox
$\mbox{S}_{dual}$(AUC) $\mbox{P}_{dual}$ $\overline{\mbox{P}}_{dual}$(AUC) $\mbox{P}_{angle}$
SiamX-360 0.205 0.278 0.278 0.327
AiATrack-360 0.362 0.449 0.516 0.535
360VOT BFoV 360VOT rBFoV
$\mbox{S}_{sphere}$(AUC) $\mbox{P}_{angle}$ $\mbox{S}_{sphere}$(AUC) $\mbox{P}_{angle}$
SiamX-360 0.262 0.327 0.243 0.323
AiATrack-360 0.548 0.564 0.426 0.530
BBox
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
rBBox
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
BFoV
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
rBFoV
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Figure 9: Qualitative tracking results of the baseline on different representations of 360VOT. Red denotes the ground truth, and blue denotes the results of AiATrack-360. The green in the first row denotes the results of AiATrack.
Overall performance on BBox. Existing trackers take the BBox of the first frame to initialize the tracking, and the inference results are also in the form of BBox. Table III shows comparison results among four groups of trackers, i.e., other, Siamese, transformer baselines, and the adapted trackers for each block in the table. According to the quantitative results, PrDiMP [57], SiamX [37] and AiATrack-360 perform best in their group of trackers. Owing to the powerful network architectures, the transformer trackers generally outperform other groups of the compared trackers. After AiATrack integrates our proposed framework, AiATrack-360 achieves a significant performance increase of 12.9%, 13.7%, 13.6% and 15.1% in terms of Sdual, Pdual, $\overline{\mbox{P}}_{dual}$ and Pangle respectively. AiATrack-360 outperforms all other trackers with a big performance gap. Compared to SiamX, SiamX-360 is improved by 8.9% Sdual, 10% Pdual, 9.6% $\overline{\mbox{P}}_{dual}$ and 11% Pangle, which is comparable with other transformer trackers. Although the performance gains of AiATrack-360 and SiamX-360 are different, the difference validates the effectiveness and generalization of our 360 tracking framework on 360°$\degree$ visual object tracking. They can serve as a new baseline for future comparison.
Attribute-based performance on BBox. Furthermore, we evaluated all trackers under 20 attributes to analyze different challenges. In Figure 8, we plot the results on the videos with cross border (CB), fast motion on the sphere (FMS), and latitude variation (LV) attributes. These are the three exclusive and most challenging attributes of 360VOT. Compared to the overall performance, all trackers suffer performance degradation, especially on the CB and FMS attributes. For example, Pdual of SimTrack decreases by 4.2% and 5.3% on CB and FMS respectively. However, the performance of AiATrack-360 still dominates on all three adverse attributes, while SiamX-360 also obtains stable performance gains.
Performance based on other annotations. Apart from BBox, we provide additional ground truth, including rBBox and (r)BFoV. As our 360 tracking framework can estimate approximate rBBox and (r)BFoV from local BBox predictions, we additionally evaluate the performances of SiamX-360 and AiATrack-360 on these three representations (Table IV). Compared with the results on BBox (Table III), the performance on rBBox declines vastly. SiamX-360 and AiATrack-360 only achieve 0.205 and 0.362 Sdual respectively. By contrast, the evaluation of (r)BFoV has more reasonable and consistent numbers. In addition, we display visual results by AiATrack-360 and AiAtrack in Figure 9. AiATrack-360 can follow and localize targets in challenging cases. Compared with (r)BBox, (r)BFoV can bind targets accurately with fewer irrelevant areas. From the extensive evaluation, we observe that (r)BFoV is beneficial for object localization in omnidirectional scenes.
V-C VOS Evaluation
VOS baseline trackers. On 360VOS benchmark, we tested 16 state-of-the-art video object segmentation trackers. Among them, we included five memory-based methods that have demonstrated impressive results in VOS tasks: STM[6], GMVOS[58], STCN[59], XMem[10], and XMem++[60]. Additionally, we evaluated unified VOS methods across segmentation and tracking namely RTS[38], UNICORN[7], and TarViS[61], transformer-based methods including AOT[62] and DeAOT[9], and other learning-based methods including AFB-URR[63], CFBI[8], CFBI+[64], LWL[65], TBD[66], and JOINT[67]. In the experiments, some listed trackers had excessive GPU and RAM demands when handling a long sequence (i.e., 2400 frames) of 360VOS. For memory-based trackers STCN[59], GMVOS[58], and STM[6], we limited their memory clip length to 16. As for trackers AOT[62] and DeAOT[9] that integrate long-term memory, we only kept the latest 32 clips of their long-term memory. Furthermore, we combined XMem[10] with our 360 tracking framework as a framework baseline XMem-360. To verify the efficacy of the 360VOS training set, we retrained XMem [10] to obtain a retrained baseline XMem and evaluated it without our framework. Finally, we applied the retrained model to the 360 tracking framework, denoted as XMem-360.
TABLE V: Segmentation performance on 360VOS in terms of standard and spherical region similarity and contour accuracy. Bold blue indicates the best results in existing trackers, while bold red indicates the best results overall.
VOS Tracker 360VOS
[…]
STM[6] 0.364 0.438 0.351 0.426
UNICORN[7] 0.339 0.469 0.328 0.461
TarVis[61] 0.325 0.411 0.320 0.407
AFB-URR[63] 0.389 0.481 0.375 0.472
GMVOS[58] 0.431 0.522 0.416 0.508
CFBI[8] 0.412 0.513 0.397 0.502
CFBI+[64] 0.428 0.532 0.414 0.525
DeAOT[9] 0.447 0.578 0.433 0.569
AOT[62] 0.471 0.597 0.457 0.589
LWL[65] 0.469 0.567 0.457 0.559
TBD[66] 0.474 0.598 0.459 0.590
JOINT[67] 0.487 0.587 0.476 0.581
RTS[38] 0.540 0.645 0.530 0.641
STCN[59] 0.550 0.667 0.535 0.657
XMem[10] 0.560 0.662 0.550 0.659
XMem++[60] 0.579 0.692 0.570 0.687
Xmem 0.596 0.703 0.584 0.699
XMem-360 0.656 0.780 0.639 0.771
XMem-360 0.665 0.786 0.651 0.779
Training details of XMem. We adhered to a similar training scheme of the original XMem[10] to retrain the weights of the XMem network. We first utilized its initialized weights that were pre-trained on static images and BL30K dataset[68], and then we directly performed the main training stage on the combination of our 360VOS training set, YouTubeVOS[19] and DAVIS[18]. To reduce memory cost while learning details on the high-resolution 360°$\degree$ images of 360VOS, we downsampled each frame from 4K resolution to 720p (i.e., 1440$\times$720) for training. During training, all training frames were randomly cropped to $512\times 512$. We employed bootstrapped cross entropy loss and dice loss functions to optimize the model. The optimizer we used is AdamW while the learning rate is set to 1e-5 with a weight decay rate of 0.05. The retraining process lasted around 3.5 days, involving 100K iterations with a batch size of 8 on 4 RTX3090 GPUs.
GT
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
XMem
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
XMem
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
XMem-360
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
XMem-360
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Figure 10: Comparison between the ground truth (GT) and the results from XMem[10] and our three new baselines. The output masks and BFoVs are shown with distinct colors. The 360 tracking framework enhances performance in addressing specific challenges, such as crossing border (CB) and large distortion (LD), while the retraining strategy (i.e. XMem*, XMem-360*) yields noticeable improvements on large distortion and stitching artifact (SA).
Performance on 360VOS. In each testing sequence, the mask was initialized on the first frame, and all trackers subsequently inferred masks for the following frames. According to the quantitative results reported in Table V, memory-based methods such as STCN[59], XMem[10], and XMem++[60] outperform other existing trackers when dealing with omnidirectional scenes. Particularly, XMem++ [60] achieve the top performance among the 16 existing trackers, with scores of 0.579, 0.692, 0.570, and 0.687 in terms of $\mathcal{J}$, $\mathcal{F}$, $\mathcal{J}_{\text{sphere}}$, and $\mathcal{F}_{\text{sphere}}$, respectively. By leveraging the 360VOS data knowledge, the retrained baseline XMem still show a noticeable improvement of 1.7%, 1.1%, 1.4%, and 1.2% compared to XMem++[60] in terms of $\mathcal{J}$, $\mathcal{F}$, $\mathcal{J}_{\text{sphere}}$, and $\mathcal{F}_{\text{sphere}}$. With our proposed framework, both XMem-360 and XMem-360 demonstrate significant improvements over all other trackers. Notably, XMem-360 and XMem achieve slightly better performance than XMem-360 and XMem, respectively, which present the effectiveness of the 360VOS training set in addressing the intricate appearance variance of tracking targets in 360VOS tasks, as exemplified in Figure 10.
TABLE VI: Tracking performance of adapting VOS trackers on the 360VOT benchmark. (VOT) indicates the (r)BBox and (r)BFoV results of the trackers are converted from their original output masks. In the same group of performance, bold red highlights the best results, while bold blue highlights the second results.
Tracker 360VOT BBox
$\mbox{S}_{dual}$(AUC) $\mbox{P}_{dual}$ $\overline{\mbox{P}}_{dual}$(AUC) $\mbox{P}_{angle}$
SiamX-360 0.391 0.365 0.430 0.425
AiATrack-360 0.534 0.506 0.563 0.574
(VOT) XMem[10] 0.432 0.425 0.460 0.465
(VOT) XMem 0.523 0.529 0.553 0.576
(VOT) XMem-360 0.535 0.534 0.566 0.588
(VOT) XMem-360 0.583 0.588 0.607 0.637
Tracker 360VOT rBBox
$\mbox{S}_{dual}$(AUC) $\mbox{P}_{dual}$ $\overline{\mbox{P}}_{dual}$(AUC) $\mbox{P}_{angle}$
SiamX-360 0.205 0.278 0.278 0.327
AiATrack-360 0.362 0.449 0.516 0.535
(VOT) XMem[10] 0.288 0.411 0.430 0.457
(VOT) XMem 0.347 0.511 0.517 0.564
(VOT) XMem-360 0.357 0.512 0.524 0.574
(VOT) XMem-360 0.387 0.566 0.565 0.624
Tracker 360VOT BFoV 360VOT rBFoV
$\mbox{S}_{sphere}$(AUC) $\mbox{P}_{angle}$ $\mbox{S}_{sphere}$(AUC) $\mbox{P}_{angle}$
SiamX-360 0.262 0.327 0.243 0.323
AiATrack-360 0.548 0.564 0.426 0.530
(VOT) XMem[10] 0.435 0.465 0.414 0.458
(VOT) XMem 0.523 0.575 0.498 0.568
(VOT) XMem-360 0.541 0.585 0.514 0.577
(VOT) XMem-360 0.594 0.635 0.565 0.626
Tracking performance of VOS trackers on 360VOT. To explore the potential for cross-domain methodological enhancements that integrate the strength of VOS and VOT methods, we conducted additional experiments to evaluate the tracking performance of VOS trackers on the 360VOT benchmark. In the implementation, we converted the VOS outputs of XMem[10] and the 3 new baselines to obtain the tracking results in terms of (r)BBox and (r)BFoV. The comparisons with the pure VOT trackers (i.e., AiATrack-360 and SiamX-360) are demonstrated in Table VI. XMem[10] exhibits commendable performance in VOT tasks surpassing SiamX-360. Compared to AiATrack-360, XMem-360 shows comparable performance while the tracking precision is higher. For example, the rBBox angle precision Pangle of XMem-360 outperforms AiATrack-360 by 3.9%. It indicates an outstanding capability of VOS trackers in handling objects with complex orientations within 360°$\degree$ videos. This is further evidenced by an obvious performance gain in the rBFoV evaluation. Moreover, XMem-360 achieves consistent advancements across all tracking metrics, proving the effectiveness of the training set.
VI Discussion and Conclusion
In this paper, we explore omnidirectional visual object tracking and segmentation. We take advantage of BFoV to represent target positions on 360°$\degree$ images and extract less-distorted searching regions. The BFoV representation thus becomes the wheel in our 360 tracking framework that can adopt arbitrary classical local visual trackers to perform visual object tracking or segmentation in omnidirectional videos. The integration of existing algorithms into our 360 tracking framework demonstrates promising results in handling the intricate challenges posed by spherical distortions on 360°$\degree$ images. In addition, we introduced a comprehensive benchmark dataset 360VOTS designed to facilitate the study of both omnidirectional visual object tracking and segmentation. The newly available training data also provides an effective way to enhance the tracker’s performance.
Nevertheless, there still remains a big room for improvement and we want to discuss some promising directions here. One of them is the development of long-term omnidirectional tracking algorithms. The trackers enhanced by our tracking framework are still classified as short-term trackers technically. As target occlusion is a noticeable attribute of 360VOTS, the long-term tracker capable of target relocalization can perform better. However, how to effectively and efficiently search for targets on the whole 360°$\degree$ image is a challenge. Another promising direction is exploring new network architectures. The trackers exploiting network architectures [69, 70] tailored for omnidirectional images may be able to extract better features and correlations for robust VOT and VOS.
By introducing new representations, new datasets, new metrics, and new benchmarks, we envisage 360VOTS will not only catalyze further research on omnidirectional visual object tracking and segmentation but also promote new applications across a broad spectrum of computer vision and robotics fields.